2.) Find the limit or show it does not exist V1 + 4x6 lim x-- 2 - 83
Answer #1

+4 6 x V1+4x lim x+40 2-X lim 2 2 73 +4 = lim +4 = lim -1 x +4 = = lim 2 10+4 (0-1) 2 -1 = -2